FIND THE EXPECTED VALUE USING ADDITION RULES:
A 35 year old woman purchases a $100,000 term life insurance policy for an annual payment of $360. Based on a period life table for the U.S. government , the probability that she will survive the year is 0.999057. Find the expected value of the policy for the insurance company.
could you show me how this is done.
1
Expert's answer
2012-04-24T07:18:18-0400
The expected payout for each of the insured in a given year:

1-0.999057 = 0.000943*100.000 = $94.3

Thus, their profit for the first year is

$360 - $94.3 = $265.7.